OpenStudy (anonymous):

What operation must be done to both sides of the equation to solve p minus negative three and seven-ninths equals five-sixths?

OpenStudy (anonymous):

its Add three and seven-ninths

jimthompson5910 (jim_thompson5910):

p minus negative three and seven-ninths equals five-sixths translates to \[\Large p - \left(-3\frac{7}{9}\right) = \frac{5}{6}\]

jimthompson5910 (jim_thompson5910):

Which simplifies to \[\Large p + 3\frac{7}{9} = \frac{5}{6}\]